# Hierarchical Multi-agent Reinforcement Learning for Cyber Network Defense

### v2

@@ Abstract.
Recent advances in multi-agent reinforcement learning (MARL) have created opportunities to solve complex real-world tasks. Cybersecurity is a notable application area, where defending networks against sophisticated adversaries remains a challenging task typically performed by teams of security operators. In this work, we explore novel MARL strategies for building autonomous cyber network defenses that address challenges such as large policy spaces, partial observability, and stealthy, deceptive adversarial strategies. To facilitate efficient and generalized learning, we propose a hierarchical Proximal Policy Optimization (PPO) architecture that decomposes the cyber defense task into specific sub-tasks like network investigation and host recovery. Our approach involves training sub-policies for each sub-task using PPO enhanced with domain expertise. These sub-policies are then leveraged by a master defense policy that coordinates their selection to solve complex network defense tasks. Furthermore, the sub-policies can be fine-tuned and transferred with minimal cost to defend against shifts in adversarial behavior or changes in network settings. We conduct extensive experiments using CybORG Cage 4, the state-of-the-art MARL environment for cyber defense. Comparisons with multiple baselines across different adversaries show that our hierarchical learning approach achieves top performance in terms of convergence speed, episodic return, and several interpretable metrics relevant to cybersecurity, including the fraction of clean machines on the network, precision, and false positives on recoveries.
Key words and phrases:
Multi-agent reinforcement learning, Cybersecurity, Deep reinforcement learning, Hierarchical reinforcement learning
1. Introduction
@@
Cyber defense is critical in both private and public network infrastructures, which are frequently targeted by increasingly sophisticated external attackers with malicious intentions. In 2024, the number of security breaches has surpassed 10,000 and attackers constantly adapt their tools and strategies to evade existing defenses Hylender et al. (2024). The implications for national security are significant, as security breaches often lead to theft of intellectual property, compromise of sensitive information, and disruption of critical infrastructures. Currently, organizations employ teams of security professionals who constantly oversee the security of their networks and design the overall security strategy using their domain expertise. While a range of machine learning (ML) tools are available for detecting specific classes of attacks Antonakakis et al. (2012); Nelms et al. (2013); Antonakakis et al. (2011); Bilge et al. (2012); Yen et al. (2013); Ongun et al. (2021), the advancement of deep reinforcement learning (DRL) presents an opportunity to automate the cyber defense strategy and reduce the burden on security operators.
@@
Towards this goal, the technical cooperation program (TTCP), a collaborative working group including UK, USA, Canada, Australia and New Zealand, developed a series of CAGE challenges for advancing cyber defense cag (2021). These challenges leverage the Cyber Operations Research Gym (CybORG), a simulated environment that can be used for creating realistic interactions (or games) between attackers and defenders on realistic network topologies. These environments task defenders (blue agents) with monitoring and restoring compromised machines on a simulated enterprise network, to prevent external adversaries (red agents) from accessing critical assets. The first CAGE challenges model a cyber game between a single blue agent and a single red agent, leading to the development of DRL techniques for training a blue agent interacting against a red agent Vyas et al. (2023); Mcdonald et al. (2024); Hammar et al. (2024). The most recent CAGE 4 challenge Group (2023) models a team of multiple blue agents defending a distributed network, playing against multiple red agents compromising the network. Existing techniques for single defensive agents are either computationally expensive (by training a different agent for each red agent Vyas et al. (2023); Mcdonald et al. (2024)), do not generalize to new attackers, or require extensive, causal pre-processing which is intractable as the network size scales Hammar et al. (2024). Thus, they cannot be immediately applied to the multi-agent CAGE 4 environment, which requires new methods. Additional challenges for training multi-agent defenders in this environment include large policy spaces, partial observability of the network, shared rewards among all blue agents, and playing against stealthy, deceptive adversarial strategies.
@@
In this paper, we propose the first scalable multi-agent reinforcement learning (MARL) technique for automating defense in cyber security environments such as CAGE 4. We formulate the problem as a decentralized, partially-observable Markov decision process (Dec-POMDP) Oliehoek and Amato (2016) and introduce a novel hierarchical multi-agent training algorithm utilizing security domain expertise. Our method decomposes the complex cyber defense task into smaller sub-tasks, trains sub-policies for each sub-task using PPO enhanced with domain expertise, and finally trains a master policy that coordinates the selection of the sub-policies at each time step. Another insight in our design is that extended observation spaces including security indicators, such as presence of malicious files and malicious processes on a host, are beneficial in increasing the blue agent’s ability to defend the network. We evaluate our method against multiple baselines at different stages of the design process to motivate our methodology and design decisions. We further propose multiple relevant and interpretable metrics for cyber network defense, including ratio of uninfected hosts, false/true positive rates on host recovery, and number of adversarial impacts on hosts. Across these metrics our proposed hierarchical techniques display significant improvements over naive MARL approaches.
@@
To summarize, our contributions are: (1) We propose scalable hierarchical multi-agent reinforcement learning methods for cyber defense; (2) We guide our design using domain expertise to enhance the agents’ observation space and decompose the complex cyber defense task into multiple sub-tasks; (3) We evaluate our methods in CybORG CAGE 4, a realistic cyber environment with partial observability and deceptive, stealthy adversaries; (4) We show transferability of trained sub-policies after fine-tuning to new adversarial agents, and (5) introduce multiple interpretable metrics for providing insights to security operators. The source code will be made publicly available on GitHub.
2. Related Work and Background
@@
Traditional cyber defenses, such as anti-virus and network intrusion detection tools, leverage specific detection rules for thwarting existing attacks, but they are relatively easy to evade. To address their limitations, organizations employ security operators who perform “threat hunting” to detect novel attacks on their networks. A variety of machine learning (ML) tools are available for threat detection Antonakakis et al. (2011, 2012); Bilge et al. (2012); Nelms et al. (2013); Yen et al. (2013); Ongun et al. (2021), but the overall defensive strategy in most organizations is still manually designed. The advancement of DRL and MARL provides an opportunity to automate cyber defense strategies and improve the security of cyber infrastructures.
@@
The CAGE-4 challenge Group (2023) is a recent security environment aimed at encouraging research in autonomous cyber defense. It provides a cyber simulation of attacker and defender actions in realistic network topologies. CAGE-4 is a partially observable environment with multiple, decentralized blue agents defending the network by playing against a team of red agents performing various attacks over time. It can be modeled as a decentralized, partially observable Markov decision process (Dec-POMDP). Dec-POMDPs Oliehoek and Amato (2016) are a special class of MDP where multiple, independent and decentralized agents with incomplete observations interact to optimize a shared reward signal. Formally, a Dec-POMDPs is defined as the tuple $\mathcal{M}=(\mathcal{I},\mathcal{S},\mathcal{A},\mathcal{T},\Omega,\mathcal{O},\mathcal{R},b_{o})$, where =$\mathcal{I}=\{1\cdots n\}$ is the set of $n$ agents, $\mathcal{S}$ is a finite set of states, $\mathcal{A}=\times_{i\in\mathcal{I}}A_{i}$ is the set of joint actions composed of individual actions $A_{i}$ for each agent $i$, $\mathcal{T}$ is the state transition function, $\Omega=\times_{i\in\mathcal{I}}\Omega_{i}$ is the set of joint observations, $\mathcal{O}:\mathcal{S}\times\mathcal{A}\rightarrow\Omega$ is the joint observation function, and $b_{0}$ is a distribution over initial states.
Several methods have been developed for solving general Dec-POMDPs, including multi-agent PPO Yu et al. (2022); Schulman et al. (2017), Q-MIX Rashid et al. (2020), and independent PPO De Witt et al. (2020). However, these methods are often uninterpretable and struggle to converge in settings with large joint action spaces. Consequently, some works have developed hierarchical or meta-action-based approaches, which combine multiple actions or sequences of actions into a single, meta-action.
Hierarchical and meta-learning approaches in single-agent reinforcement learning have been explored Hutsebaut-Buysse et al. (2022); Lake et al. (2017); Hospedales et al. (2022); Mishra et al. (2018); Sun et al. (2019); Guan et al. (2024). Several works have developed adaptations of hierarchical approaches in MARL for different domains Xiao et al. (2020); Fosong et al. (2024); Chang et al. (2023); Frans et al. (2017), though the general problem remains a challenge. Therefore, in this work, we aim to study the design and capabilities of Hierarchical MARL (H-MARL) approaches specific to the cyber network defense domain. Previous works Vyas et al. (2023); Mcdonald et al. (2024); Hammar et al. (2024) have looked at single-agent domains, and to the best of our knowledge there is no existing work applying MARL to cyber network defense.
3. Problem Statement
@@
Table 1. Challenges of a realistic RL model for cybersecurity, with concrete examples from the CybORG environment.
@@
In cyber security, red teams act as attackers who attempt to exploit network vulnerabilities and carry out malicious activities aimed at compromising the system. Blue teams are tasked with defending against red team opponents to secure the networks, while maintaining network operations. In this work, we focus on the CybORG CAGE 4 cybersecurity MARL framework Group (2023), which is a realistic environment that models cyber defense. We discuss several aspects of the environment below. For additional details, please see Section A.1 in the supplemental materials, and the CAGE 4 description Group (2023).
Threat Model. The two teams are represented by multi-agent systems: defender (the blue team) and attacker (the red team). Defender and attacker have competing goals, while the agents on each team collaborate to achieve their goals. The attacker’s goal is to maximize its reward by degrading services available to users, represented by green agents, and compromising the critical Operational Technology (OT) service. The defender’s goals are two-fold: maintain the security of the cyber network by reducing the adversarial presence, and minimize the operational impact on users.
We face a strong adversary, whose capabilities include stealth, phishing, propagating through the network, and the ability to discover blue agents’ deception. Red agents maintain persistent presence on the Contractor network, which is the starting point of the attack and cannot be defended with blue actions. Red agents scan machines for vulnerable services to exploit and propagate through the network. CAGE 4 implements both an aggressive service discovery action, which is faster (1 time step), but has a high chance (0.75) of raising alerts, and a stealthy service discovery action, which is slower (3 time steps), but less likely (0.25) to raise alerts. In addition to moving laterally through the network by exploiting remote services, red agents can also spawn with a given probability when a green user opens a phishing email or accesses a compromised service. Furthermore, red agents can use the discover deception action to determine if the blue team has installed decoy services on a specific host, and avoid to infect that host to maintain stealth. CAGE 4 implements a default deceptive red agent, called FiniteStateRedAgent, but we also create our own red agents with more aggressive service discovery and even stealthier presence on the network to measure the generality of our defense.
@@
Rewards. The reward scheme models a zero-sum game where blue agents incur penalties when green agents are impacted due to degraded services becoming inaccessible. In addition, blue agents are penalized when red agents impact the critical OT security service. The specific reward values depend on the mission phase. Three active phases are carried out throughout each episode, to reflect the changing criticality of security zones on current operations.
@@ 4. H-MARL Methodology
For complex real-world tasks, action spaces can grow intractably large as the problem, or task, scales, making task learning difficult for many standard reinforcement learning (RL) approaches Dulac-Arnold et al. (2015). These issues are compounded by the introduction of high-dimensional, noisy state spaces. Many hierarchical RL methods aim to solve these problem by breaking large action spaces down into smaller sub-tasks, however learning these partitions online remains a challenging problem Hutsebaut-Buysse et al. (2022). Therefore, in this section, we introduce our hierarchical framework demonstrating how domain expertise can be leveraged to solve issues of intractability in both the state and action spaces.
@@
Figure 1. Hierarchical MARL. Upon receiving an observation, the master policy first chooses a sub-policy, which selects the final primitive action.
@@
In our proposed design, shown in Figure 1, the action space $\mathcal{A}$ is first broken into a smaller set of $k$ meta-actions $\mathcal{A}_{m}=\{A_{c}\}_{c=1}^{k}$ where each $A_{c}\subset\mathcal{A}$ is a disjoint subset of actions chosen using domain expertise. In cyber domains the agent must chose which host machine, potentially from a list of thousands, to investigate or restore. Under our design, these action will instead be wrapped into a single meta-action, significantly reducing the action space. Under this formulation we begin to leverage our hierarchical design. We first establish a master policy $\pi_{m}$ whose objective is to choose some meta-action $A_{c}\in\mathcal{A}_{m}$ given observation $o_{t}$ at time step $t$.
@@
We then assign each meta-action $A_{c}$ to a corresponding sub-policy $\psi_{c}$ whose goal is to choose the primitive action $a_{t}\in A_{c}$ given some input history $h_{t}$ at time step $t$. This primitive action $a_{t}$ is the final action executed in the environment. Under this design the master policy must learn the best policy $\pi_{m}:\mathcal{H}\rightarrow\mathcal{A}_{m}$ over meta-actions, while each sub-policy $\psi_{c}:\mathcal{H}\rightarrow A_{c}$ must learn the best policy over all actions in their respective class of meta-action. Here $\mathcal{H}$ represents the set of possible observation histories. This reduces the larger, more complex task posed by the base Dec-POMDP into a much more manageable set of sub-tasks.
We then use expert domain knowledge to enhance each sub-policy’s respective observations with transformation functions $f_{c}:\mathcal{O}\rightarrow\mathcal{O}_{c}$ for sub-policy observation spaces $\mathcal{O}_{c}$. In practice, these transformations are applied to observation histories, which will be explained in Section 4.2. These functions make learning easier for each sub-policy by removing information irrelevant to their respective class of meta-action. In cyber domains, it is important for a sub-policy responsible with monitoring the network to observe all suspicious events, allowing them to focus their investigations. On the other hand, a sub-policy responsible for recovering machines from compromised states would be better off only observing indicators of compromise (IoCs) which are more direct indicators that the adversary has gained root or host-level access. In the following subsections, we explore this design in more detail.
4.1. Hierarchical MARL Design

### v3

@@ Abstract
Recent advances in multi-agent reinforcement learning (MARL) have created opportunities to solve complex real-world tasks. Cybersecurity is a notable application area, where defending networks against sophisticated adversaries remains a challenging task typically performed by teams of security operators. In this work, we explore novel MARL strategies for building autonomous cyber network defenses that address challenges such as large policy spaces, partial observability, and stealthy, deceptive adversarial strategies. To facilitate efficient and generalized learning, we propose a hierarchical Proximal Policy Optimization (PPO) architecture that decomposes the cyber defense task into specific sub-tasks like network investigation and host recovery. Our approach involves training sub-policies for each sub-task using PPO enhanced with cybersecurity domain expertise. These sub-policies are then leveraged by a master defense policy that coordinates their selection to solve complex network defense tasks. Furthermore, the sub-policies can be fine-tuned and transferred with minimal cost to defend against shifts in adversarial behavior or changes in network settings. We conduct extensive experiments using CybORG Cage 4, the state-of-the-art MARL environment for cyber defense. Comparisons with multiple baselines across different adversaries show that our hierarchical learning approach achieves top performance in terms of convergence speed, episodic return, and several interpretable metrics relevant to cybersecurity, including the fraction of clean machines on the network, precision, and false positives.
1 Introduction
@@
Cyber defense is critical in both private and public network infrastructures, which are frequently targeted by increasingly sophisticated external attackers with malicious intentions. In 2024, the number of security breaches has surpassed 10,000 and attackers constantly adapt their tools and strategies to evade existing defenses (Verizon, ). The implications for national security are significant, as security breaches often lead to theft of intellectual property, compromise of sensitive information, and disruption of critical infrastructures. Currently, organizations employ teams of security professionals who constantly oversee the security of their networks and design the overall security strategy using their domain expertise. While a range of machine learning (ML) tools are available for detecting specific classes of attacks (180232, ; nelms2013execscent, ; antonakakis11, ; EXPOSURE, ; yen2013beehive, ; portfiler, ), the advancement of deep reinforcement learning (DRL) presents an opportunity to automate the cyber defense strategy and reduce the burden on security operators.
@@
Towards this goal, the technical cooperation program (TTCP), a collaborative working group including UK, USA, Canada, Australia and New Zealand, developed a series of CAGE challenges for advancing cyber defense (cage_challenge_announcement, ). These challenges leverage the Cyber Operations Research Gym (CybORG), a simulated environment that can be used for creating realistic interactions (or games) between attackers and defenders on realistic network topologies. These environments task defenders (blue agents) with monitoring and restoring compromised machines on a simulated enterprise network, to prevent external adversaries (red agents) from accessing critical assets. The first CAGE challenges model a cyber game between a single blue agent and a single red agent, leading to the development of DRL techniques for training a blue agent interacting against a red agent (vyas2023automated, ; 10639381, ; hammar2024optimal, ). The most recent CAGE 4 challenge (cage_challenge_4_announcement, ) models a team of multiple blue agents defending a distributed network, playing against multiple red agents compromising the network. Existing techniques for single defensive agents are either computationally expensive (by training a different agent for each red agent (vyas2023automated, ; 10639381, )), do not generalize to new attackers, or require extensive, causal pre-processing which is intractable as the network size scales (hammar2024optimal, ). Thus, they cannot be immediately applied to the multi-agent CAGE 4 environment, which requires new methods. Additional challenges for training multi-agent defenders in this environment include large policy spaces, partial observability of the network, shared rewards among all blue agents, and playing against stealthy, deceptive adversarial strategies.
@@
In this paper, we propose the first scalable multi-agent reinforcement learning (MARL) technique for automating defense in cyber security environments such as CAGE 4. We formulate the problem as a decentralized, partially-observable Markov decision process (Dec-POMDP) (Oliehoek2016ACI, ) and propose two hierarchical strategies, H-MARL Expert and H-MARL Meta, each with their own advantages. Both methods decompose the complex cyber defense task into smaller sub-tasks, and train sub-policies for each sub-task using PPO enhanced with domain expertise. The difference between the methods is in the design of the master policy that coordinates the selection of the sub-policies at each time step. H-MARL Expert utilizes security domain expertise to define a top master policy based on well-established practices for cyber defense, and performs best in most of our experiments. However, there are situations when it is difficult to define a deterministic Expert policy. To address this issue, we propose H-MARL Meta, that trains the master policy, and, thus, has the advantage of generalizing to new, unseen adversarial behavior. Another insight in our design is that extended observation spaces including security indicators, such as presence of malicious files and malicious processes on a host, are beneficial in increasing the blue agent’s ability to defend the network. We evaluate our methods against multiple baselines at different stages of the design process to motivate our methodology and design decisions. We further propose multiple relevant and interpretable metrics for cyber defense, including ratio of uninfected hosts, false/true positive rates on host recovery, and number of adversarial impacts on hosts. Across these metrics our proposed hierarchical techniques display significant improvements over traditional MARL approaches.
@@
To summarize, our contributions are: (1) scalable hierarchical multi-agent reinforcement learning methods for cyber defense; (2) a design guided by domain expertise to enhance the agents’ observation space and decompose the complex cyber defense task into multiple sub-tasks; (3) evaluation in CybORG CAGE 4, a realistic cyber environment with partial observability and deceptive, stealthy adversaries; (4) empirical transferability of trained sub-policies after fine-tuning to new adversarial agents, and (5) multiple interpretable metrics for providing insights to security operators. The source code is available on GitHub  [^1] https://github.com/adityavs14/Hierarchical-MARL.
2 Related Work and Background
@@
Traditional cyber defenses, such as anti-virus and network intrusion detection tools, leverage specific detection rules for thwarting existing attacks, but they are relatively easy to evade. To address their limitations, organizations employ security operators who perform “threat hunting” to detect novel attacks on their networks. A variety of machine learning (ML) tools are available for threat detection (antonakakis11, ; 180232, ; EXPOSURE, ; nelms2013execscent, ; yen2013beehive, ; portfiler, ), but the overall defensive strategy in most organizations is still manually designed. The advancement of DRL and MARL provides an opportunity to automate cyber defense strategies and improve the security of cyber infrastructures.
@@
The CAGE-4 challenge (cage_challenge_4_announcement, ) is a recent security environment aimed at encouraging research in autonomous cyber defense. It provides a cyber simulation of attacker and defender actions in realistic network topologies. CAGE-4 is a partially observable environment with multiple, decentralized blue agents defending the network by playing against a team of red agents performing various attacks over time. It can be modeled as a decentralized, partially observable Markov decision process (Dec-POMDP). Dec-POMDPs (Oliehoek2016ACI, ) are a special class of MDP where multiple, independent and decentralized agents with incomplete observations interact to optimize a shared reward signal. Formally, a Dec-POMDPs is defined as the tuple $\mathcal{M}=(\mathcal{I},\mathcal{S},\mathcal{A},\mathcal{T},\Omega,\mathcal{O},\mathcal{R},b_{o})$, where $\mathcal{I}=\{1\cdots n\}$ is the set of $n$ agents, $\mathcal{S}$ is a finite set of states, $\mathcal{A}=\times_{i\in\mathcal{I}}A_{i}$ is the set of joint actions composed of individual actions $A_{i}$ for each agent $i$, $\mathcal{T}$ is the state transition function, $\Omega=\times_{i\in\mathcal{I}}\Omega_{i}$ is the set of joint observations, $\mathcal{O}:\mathcal{S}\times\mathcal{A}\rightarrow\Omega$ is the joint observation function, and $b_{0}$ is a distribution over initial states.
Several methods have been developed for solving general Dec-POMDPs, including multi-agent PPO (MAPPO) (mappo, ; schulman2017proximal, ), Q-MIX (qmix, ), independent PPO (IPPO) (de2020independent, ), or decision trees. However, these methods are often uninterpretable and struggle to converge in settings with large joint action spaces. These approaches have usually been applied to simpler 2-player environments (vyas2023automated, ; 10639381, ; hammar2024optimal, ; wilson2024marlcyber, ). Prior works in the more complex CAGE-4 environment explore defenses based on heuristics (Kiely2025, ) or on traditional PPO algorithms (wang2024marlcyber, ). In contrast, we implement an observation-enhanced hierarchical learning method that is more adaptive than previous heuristic approaches, and more scalable than single-policy PPO architectures.
Hierarchical and meta-learning methods in reinforcement learning have led to adaptations of hierarchical MARL for different domains, such as multi-robot teamwork tasks (9196684, ; Fosong24, ; chang2023hierarchical, ) and complex navigation (Frans2018, ). To the best of our knowledge, our work is the first to study the design and capabilities of hierarchical MARL approaches in the cyber defense domain.
3 Problem Statement
In cyber security, red teams act as attackers who attempt to exploit network vulnerabilities and carry out malicious activities aimed at compromising the system. Blue teams are tasked with defending against red team opponents to secure the networks, while maintaining network operations. In this work, we focus on the CybORG CAGE 4 cybersecurity MARL framework (cage_challenge_4_announcement, ), which is a realistic environment that models cyber defense. We discuss several aspects of the environment below. For additional details, please see Section 1 in the supplemental materials, and the CAGE 4 description (cage_challenge_4_announcement, ).
Threat Model. The two teams are represented by multi-agent systems: defender (the blue team) and attacker (the red team). Defender and attacker have competing goals, while the agents on each team collaborate to achieve their goals. The attacker’s goal is to maximize its reward by degrading services available to users, represented by green agents, and compromising the critical Operational Technology (OT) service. The defender’s goals are two-fold: maintain the security of the cyber network by reducing the adversarial presence, and minimize the operational impact on users. We face a strong adversary, whose capabilities include stealth, phishing, propagating through the network, and the ability to discover blue agents’ deception. Red agents maintain persistent presence on the Contractor network, which is the starting point of the attack and cannot be defended with blue actions. Red agents scan machines for vulnerable services to exploit and propagate through the network. CAGE 4 implements both an aggressive service discovery action, which is faster (1 time step), but has a high chance (0.75) of raising alerts, and a stealthy service discovery action, which is slower (3 time steps), but less likely (0.25) to raise alerts. In addition to moving laterally through the network by exploiting remote services, red agents can also spawn with a given probability when a green user opens a phishing email or accesses a compromised service. Furthermore, red agents can use the discover deception action to determine if the blue team has installed decoy services on a specific host, and avoid to infect that host to maintain stealth. CAGE 4 implements a default deceptive red agent, called FiniteStateRedAgent, but we also create our own red agents with more aggressive service discovery and even stealthier presence on the network to measure the generality of our defense.
@@ 3 Problem Statement
Table 1: Challenges of a realistic RL model for cybersecurity, with concrete examples from the CybORG environment.
@@
Rewards. The reward scheme models a general-sum game where blue agents incur penalties when green agents are impacted due to degraded services becoming inaccessible. In addition, blue agents are penalized when red agents impact the critical OT security service, or when they use a costly action like Restore machine. The specific reward values depend on the mission phase and are specified on the challenge page cage_challenge_4_announcement . Three mission phases are carried out throughout each episode, to reflect the changing criticality of security zones on current operations. Note that the reward includes the penalties incurred inside the contractor subnet, which cannot be defended. This additional reward should not affect the training process. For a fair comparison, the contractor reward is present in all the methods studied in this paper, including the baselines.
@@ 4 H-MARL Methodology
For complex real-world tasks, action spaces can grow intractably large as the problem scales, making task learning difficult for many standard reinforcement learning (RL) approaches (dulac2015deep, ). These issues are compounded by the introduction of high-dimensional, noisy state spaces. Many hierarchical RL methods aim to solve these problems by breaking large action spaces down into smaller sub-tasks. However, learning these partitions online remains a challenging problem (Hutsebaut-Buysse2022, ). Therefore, in this section, we introduce our hierarchical framework showing how domain expertise can be leveraged to solve issues of intractability in both the state and action spaces.
@@
Figure 1: Hierarchical MARL. Upon receiving an observation, the master policy first chooses a sub-policy, which selects the final primitive action.
@@
An overview of the hierarchical design is shown in Figure 1. The action space $\mathcal{A}$ is split into $n$ smaller subsets, or classes, chosen using domain expertise. For example, the “recover” class refers to all primitive actions for removing processes and restoring machines on the network. Thus, each sub-policy handles one class of primitive actions that will be executed in the network. We define the action space of the master policy as a new set $\mathcal{A}_{m}$ comprised of meta-actions, where each meta-action corresponds to a different sub-policy. The master policy chooses a meta-action, and the associated sub-policy then samples a primitive action from its subset. In cyber domains, the agent must choose the machine to investigate or restore potentially from a list of thousands. In our design, these additional details, such as what machine to restore, will be abstracted away, under a single meta-action, significantly reducing the action space.
@@
Under this formulation we first establish a master policy $\pi_{m}$ whose objective is to choose some meta-action $A_{c}\in\mathcal{A}_{m}$ given observation $o_{t}$ at time step $t$. We then assign each meta-action $A_{c}$ to a corresponding sub-policy $\psi_{c}$ whose goal is to choose the primitive action $a_{t}\in A_{c}$ given some input history $h_{t}$ at time step $t$. This primitive action $a_{t}$ is the final action executed in the environment. Under this design the master policy must learn the best policy $\pi_{m}:\mathcal{H}\rightarrow\mathcal{A}_{m}$ over meta-actions, while each sub-policy $\psi_{c}:\mathcal{H}\rightarrow A_{c}$ must learn the best policy over all actions in their respective meta-action class. Here $\mathcal{H}$ represents the set of possible observation histories. This reduces the larger, more complex task posed by the base Dec-POMDP into a much manageable set of sub-tasks.
We enhance each sub-policy’s respective observations with transformation functions $f_{c}:\mathcal{O}\rightarrow\mathcal{O}_{c}$ for sub-policy observation spaces $\mathcal{O}_{c}$. In practice, these transformations are applied to observation histories. The transformation function reduces the observation space of each sub-policy by keeping only information relevant to their respective class of actions. For example, the sub-policy responsible for restoring machines only needs to know about the hosts that present clear indicators of compromise, rather than about all the alerts in the system.
4.1 Hierarchical MARL Design